OpenStudy (anonymous):

Calculate the molarity of a solution of ethanol in water in which the mole fraction of ethanol is 0.040(assume the density of water to be one).

OpenStudy (anonymous):

Mole fraction = 0.04/1 = 1/25. In 25 moles of the total mixture, 1 mole of ethanol and 24 moles of water are present. 24 moles of water weigh 24 x 18 = 432 g = 432 mL (18g is the molar mass of water) 1 mole of ethanol is present in 432 mL of water. Molarity is no. of moles of solute present in 1 L of solution. So molarity = 1 x 1000 / 432 = 1000/432 = 2.315 M (approx.)

OpenStudy (anonymous):

but wont ethanol also occupy sum volume?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

*whistling*

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

idk meritnation says ur correct they used the same method

OpenStudy (anonymous):

molarity M= (WB/WA)*1000/V mole fraction of C2H5OH= 0.040 molar wt. of ethanol= 46gm wt. of ethanol in 1 mole of soln= 46*0.04=1.84 mole fraction of water= 0.960 molar wt. of water= 18gm wt. of water in 1 mole of solution=17.28 Volume of solution= 1.84+17.28= 19.12ml (1gm=1ml) therefore, M= (1.84/17.28)*1000/0.01921 =0.00556 M

OpenStudy (anonymous):

is that...??

OpenStudy (anonymous):

wth i cant copy but meritnations answer matches ur first

OpenStudy (vincent-lyon.fr):

Quote: "therefore, M= (1.84/17.28)*1000/0.01921 =0.00556 M" I'm afraid this is wrong. Answer is [ethanol] ≈ 0.002 mol/L

OpenStudy (anonymous):

?plz xplain

OpenStudy (anonymous):

hmmmm...

OpenStudy (vincent-lyon.fr):

Instead of using formulae, I prefer working out on an numerical example. I start with 1.00 mole of mixture. What is its composition in moles? ethanol = ........ mol water = .........mol. then for each substance, work out the mass: ethanol = ........... g water = ............ g then volume (assume density of ethanol is similar to that of water, or use the actual value of 0.79): ethanol = ............. mL water = .............. mL Final volume (ignore the slight contraction that appears when you mix ethanol and water): Final volume = .............. mL Finally: [ethanol] = ................. mol/L

OpenStudy (vincent-lyon.fr):

@shruti : sorry, I realise now I typed the same answer as yours! Only your last line was not correst, but everything else was ok!

OpenStudy (anonymous):

i dont get it how can i do the first blank it contains 1 mole ethanol and how many moles of water?

OpenStudy (vincent-lyon.fr):

By definition "mole fraction of ethanol is 0.040" means there is 0.040 mole of ethanol in 1.000 mole of mixture.

OpenStudy (anonymous):

i dont think we can assume density of eth and water to be the same. and density of ethanol would have been provided if needed...

OpenStudy (vincent-lyon.fr):

If ethanol's density is not provided, it means you have to approximate the final volume. There are 2 ways out: - ignore the volume of ethanol completely and the answer is 0.00222 - equate density of eth. and water and answer is 0.00209 Using d = 0.79 yields 0.00204 Real answer is between 0.00204 and 0.00209

OpenStudy (anonymous):

its okay vincent.